OpenStudy (anonymous):

Help, will give medal What is the smallest integer that can possibly be the sum of an infinite geometric series whose first term is 9?

OpenStudy (anonymous):

@SithsAndGiggles @misty1212

OpenStudy (welshfella):

the sum to infinity of a Geometric series is a1 / (1 - r) where a1 = first term and r is common ratio ( r < 1)

OpenStudy (anonymous):

uh... you mean a_1?

OpenStudy (welshfella):

a_1 if you like - doesnt really matter in this case its 9

OpenStudy (anonymous):

ok so the sum is 1-r and actually isnt it -1<r<1

OpenStudy (welshfella):

I'm not sure how to do this to be honest one guess would be when r = 0.1 making the sum = 9/.9 = 10

OpenStudy (welshfella):

yes it should be -1 < r < 1

OpenStudy (welshfella):

the sum is a_1 / (1 -r)

OpenStudy (welshfella):

if r = -0.5 then Sum = 9 / 1-(-0.5) = 9 / 1.5 = 6 thats lower still

OpenStudy (welshfella):

what we require is the highest negative number that will divide into 9 exactly - so we have to look at values of 1 - r where r is between -1 and 0 what if r = -0.8 then sum = 9 / (1.8) = 5

OpenStudy (welshfella):

I think thats it

OpenStudy (welshfella):

I don't think there is another number between 1.8 and 2 that gives an integer when you divide it into 9